OpenStudy (anonymous):

Two angles in a pentagon measure 65(degrees) and the third angle measures 110(degrees). What is the measurement of the remaining two congruent angles?

OpenStudy (anonymous):

help!

OpenStudy (campbell_st):

Angle sum of a polygon is 180 x ( n - 2) where n is the number of sides a pentagon n = 5 so angle sum = 180 x ( 5 -2) = 540 sum of known angles is 65 + 65 + 110 = 240 then 540 - 240 = 300 2 congruent angles just halve the answer above